# Pure C*-algebras

### v1

@@ 4. Variants of divisibility
Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible $\mathrm{Cu}$-semigroup $H$ that contains $x^{\prime},x$ and such that $H$ satisfies (O5), (O7), and (FAD).
Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}_{>0}$ such that
$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k^{2}}.$
@@
Since $H$ is countably based and satisfies (O7), we may apply Theorem 2.2. Hence, for every $s\in S$ the infimum $s\wedge\infty c^{\prime}$ exists, and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive; Set $e:=d\wedge\infty c^{\prime}$. Since $x^{\prime\prime}\ll\infty c^{\prime}$, we have
@@
Let us further choose $e^{\prime\prime}$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
There exists $N\in{\mathbb{N}}$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Choose $\varepsilon>0$ small enough such that
$\left(1+\frac{1}{k^{2}}\right)\varepsilon\leq\frac{1}{N}.$
Applying (FAD) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in H$ such that
$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq\widehat{kf}+\varepsilon\widehat{e^{\prime\prime}}.$
@@
To verify that $\widehat{x^{\prime}}\leq(k+1/k)\widehat{y}$, let $\lambda\in F(H)$. Then
$\begin{split}\lambda(x^{\prime})&\leq\frac{k(k+1)+n}{n}\lambda(e^{\prime})\leq\left(1+\frac{1}{k^{2}}\right)\lambda(e^{\prime})\\ &\leq\left(1+\frac{1}{k^{2}}\right)\left(k\lambda(f)+\varepsilon\lambda(e^{\prime\prime})\right)\leq\left(k+\frac{1}{k}\right)\lambda(f)+\frac{1}{N}\lambda(e^{\prime\prime})\\ &\leq\left(k+\frac{1}{k}\right)\lambda(f)+\lambda(c^{\prime})\leq\left(k+\frac{1}{k}\right)\lambda(y),\end{split}$
@@
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element. If $x$ is functionally almost divisible, then for every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
$y\leq x,\quad(k+\tfrac{1}{2})\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq(k+1)y.$
In particular, $x$ is upper functionally almost divisible.
@@
Let $k\in{\mathbb{N}}$. Using (FAD) with $k+1$ and $\varepsilon=1/3(2k+1)(k+1)$, we obtain an element $z\in S$ such that
$(k+1)z\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq(k+1)\widehat{z}+\frac{1}{3(2k+1)(k+1)}\widehat{x}.$
Find $\delta>0$ such that
$\frac{k+1}{1-\delta}\leq(k+1)+\frac{1}{2(2k+1)},\,\,\,\text{ and }\,\,\,\frac{\varepsilon}{(1-\delta)}\leq\frac{1}{2(2k+1)(k+1)}.$
@@
$(1-\delta)\widehat{x}\ll\widehat{x}\leq(k+1)\widehat{z}+\varepsilon\widehat{x}.$
This allows us to choose $z^{\prime}\ll z$ such that
(1) $\widehat{x}\leq\frac{k+1}{(1-\delta)}\widehat{z^{\prime}}+\frac{\varepsilon}{(1-\delta)}\widehat{x}\leq\left((k+1)+\frac{1}{2(2k+1)}\right)\widehat{z^{\prime}}+\frac{1}{2(2k+1)(k+1)}\widehat{x}$
Using Lemma 2.1, we can now find $y\in S$ such that
$kz^{\prime}+y\leq x\leq(k+1)y.$
Take $\lambda\in F(S)$. If $\lambda(z^{\prime})=\infty$, we have $\lambda(y)=\lambda(x)=\infty$ and so, in particular, $(k+\tfrac{1}{2})\lambda(y)\leq\lambda(x)$.
Thus, assume that $\lambda(z^{\prime})<\infty$. Since $k\lambda(y)\leq k(\lambda(x)-k\lambda(z^{\prime}))$, using (1) at the first step, and $(k+1)z^{\prime}\leq x$ at the last step, one has:
$\begin{split}(k+\tfrac{1}{2})\lambda(y)&\leq(k+\tfrac{1}{2})\left(\left((k+1)+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)-k\lambda(z^{\prime})\right)\\ &=(k+\tfrac{1}{2})\left(\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)\\ &\leq(k+\tfrac{1}{2})\left(\tfrac{1}{k+1}\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(x)+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)=\lambda(x),\end{split}$
as desired.
The last assertion is clear. ∎
Theorem 4.7.
Let $A$ be a $C^{*}$-algebra. Assume that $A$ is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $\operatorname{Cu}(A)$ is (upper, lower) functionally almost divisible.
@@
Assume that $A$ is $n$-almost divisible. Without loss of generality, we may assume $A$ to be stable. Further, note that we only need to prove that $\operatorname{Cu}(A)$ has (FAD) and (UFAD). Indeed, any $n$-almost divisible $C^{*}$-algebra has a $(2,\omega)$-divisible Cuntz semigroup. Thus, it follows from Lemma 4.5 that $\operatorname{Cu}(A)$ has (LFAD) whenever it has (FAD).
Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$, and let $a\in A_{+}$. Set $C=\{a\}^{\prime}\cap A_{\mathcal{U}}$ and $I=\{a\}^{\perp}\cap A_{\mathcal{U}}$. Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ has (FAD). Therefore, $[1]$ has (UFAD) by Lemma 4.6.
Now choose $k\in{\mathbb{N}}_{>0}$. Applying (FAD) to $[1]\ll[1]$ with $\varepsilon=\tfrac{1}{k+1}$, we obtain an element $e\in C/I$ such that
$k[e]\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{[e]}+\tfrac{1}{k+1}\widehat{[1]}.$
Setting $\delta=\tfrac{1}{k(k+1)}$, we have $(1+\delta)\widehat{k[1]}\leq k^{2}\widehat{[e]}+\widehat{[1]}$. Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that $(s+1)k[1]\leq s(k^{2}[e]+[1])$.
Separately, we apply (UFAD) to $[1]\ll[1]$ and we obtain an element $\tilde{f}\in C/I$ such that
$[\tilde{f}]\leq[1],\quad(k+\nu)\widehat{[\tilde{f}]}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq(k+1)[\tilde{f}],$
for some $10<\nu<$ 1. Since $[1]\ll[1]$, we can find $f\in C/I$ with $[f]\ll[\tilde{f}]$ such that $[1]\leq(k+1)[f]$. Again, applying Proposition 2.3(ii), we find a positive integer $t$ such that $(t+1)k[f]\leq t[1]$.
Let $\bar{e},\bar{f}\in C$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
$k[\bar{e}]\leq[1]+z,\quad(s+1)k[1]\leq s(k^{2}[\bar{e}]+[1])+z,$

### v3

@@ 4. Variants of divisibility
Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible sub-$\mathrm{Cu}$-semigroup $H\subseteq S$ that satisfies (O5)-(O7) and contains $x^{\prime}$, $x$.
Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}$ with $n\geq 1$ such that
$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k}.$
@@
Since $H$ is countably based and satisfies (O5)-(O7), we may apply Theorem 2.2. Hence, for every $s\in H$ the infimum $s\wedge\infty c^{\prime}$ exists (in $H$), and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive. Set $e:=d\wedge\infty c^{\prime}$. Since $x^{\prime\prime}\ll\infty c^{\prime}$, we have
@@
Let us further choose $e^{\prime\prime}\in H$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
Since $e\leq\infty c^{\prime}$ and $e^{\prime}\ll e$, we may pick $N\in{\mathbb{N}}$ with $N\geq 1$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Then choose $\varepsilon>0$ small enough such that
$\left(1+\frac{1}{k}\right)\varepsilon\leq\frac{1}{N}.$
Now, working again in $S$, and applying (L’) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in S$ such that
$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq k\widehat{f}+\varepsilon\widehat{e^{\prime\prime}}.$
@@
Further, we have
$\widehat{x^{\prime}}$ $\leq\frac{k(k+1)+n}{n}\widehat{e^{\prime}}\leq\left(1+\frac{1}{k}\right)\widehat{e^{\prime}}\leq\left(1+\frac{1}{k}\right)\left(k\widehat{f}+\varepsilon\widehat{e^{\prime\prime}}\right)$
$\leq(k+1)\widehat{f}+\frac{1}{N}\widehat{e^{\prime\prime}}\leq(k+1)\widehat{f}+\widehat{c^{\prime}}\leq(k+1)\left(\widehat{f}+\widehat{c^{\prime}}\right)=(k+1)\widehat{y},$
@@
The next result shows that if a compact element satisfies condition (L’) in Lemma 4.7, then it also satisfies condition (U) in Definition 4.3.
@@
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element such that for every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
Then, for every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\delta>0$, there exists $z\in S$ such that
$z\leq x,\quad(1-\delta)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq kz.$
@@
Given $k\in{\mathbb{N}}$ with $k\geq 1$ and $\delta>0$, pick $\varepsilon>0$ such that
$1+2(k-1)\varepsilon\leq(1-\delta)^{-1}.$
Applying the assumption for $k$ and $\varepsilon$, we obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
@@
$(1-\varepsilon)\widehat{x}\ll\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
This allows us to choose $y^{\prime}\in S$ such that $y^{\prime}\ll y$ and
$(1-\varepsilon)\widehat{x}\leq k\widehat{y^{\prime}}+\varepsilon\widehat{x}.$
Applying Lemma 2.1 for $ky\leq x$ and $y^{\prime}\ll y$, we get $z\in S$ such that
$(k-1)y^{\prime}+z\leq x\leq kz.$
We now show that $z$ has the claimed properties. It is clear that $z\leq x$ and $x\leq kz$. It remains to verify that $(1-\delta)k\lambda(z)\leq\lambda(x)$ for every $\lambda\in F(S)$. This is clear if $\lambda(x)=\infty$. Let $\lambda\in F(S)$ with $\lambda(x)<\infty$. From the fact that $(1-\varepsilon)\widehat{x}\leq k\widehat{y^{\prime}}+\varepsilon\widehat{x}$, as shown above, we deduce that
$(1-2\varepsilon)\lambda(x)\leq k\lambda(y^{\prime}).$
Since $ky\leq x$, it follows that $\lambda(y^{\prime})<\infty$ and, consequently, using our choice of $\varepsilon$ in the last step,
$k\lambda(z)$ $\leq k\lambda(x)-k(k-1)\lambda(y^{\prime})\leq k\lambda(x)-(1-2\varepsilon)(k-1)\lambda(x)$
$=(1+2(k-1)\varepsilon)\lambda(x)\leq(1-\delta)^{-1}\lambda(x),$
as desired. ∎
Proposition 4.9.
If a $C^{*}$-algebra is $n$-almost divisible for some $n\in{\mathbb{N}}$, then it is functionally divisible.
@@
Let $n\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra that is $n$-almost divisible. It is easy to deduce that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible. Therefore, by Lemma 4.7, to show that $\operatorname{Cu}(A)$ is functionally divisible it suffices to verify condition (L’) in Lemma 4.7 and condition (U) in Definition 4.3.
Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon>0$. We may assume that $\varepsilon=\tfrac{1}{N}$ for some integer $N\geq 1$, and we need to find $y\in\operatorname{Cu}(A)$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,N\widehat{x^{\prime}}\leq Nk\widehat{y}+\widehat{x},$
and $z\in\operatorname{Cu}(A)$ such that
$z\leq x,\quad(N-1)k\widehat{z}\leq N\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
Without loss of generality, we may assume that $A$ is stable, which allows us to pick $a\in A_{+}$ with $x=[a]$.
Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$. Denote by $A_{\mathcal{U}}$ the free ultrapower of $A$. Set $C=\{a\}^{\prime}\cap A_{\mathcal{U}}$ and $I=\{a\}^{\perp}\cap A_{\mathcal{U}}$. Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ satisfies condition (L’) in Lemma 4.7, that is, for every $l\in{\mathbb{N}}$ with $l\geq 1$, and every $\delta>0$, there exists $v\in\operatorname{Cu}(C/I)$ such that
$lv\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq l\widehat{v}+\delta\widehat{[1]}.$
Applied for $l=k$ and $\delta=\tfrac{1}{2N}$, we obtain $v\in\operatorname{Cu}(C/I)$ such that
$kv\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{v}+\frac{1}{2N}\widehat{[1]}.$
Adding $\tfrac{1}{2N}\widehat{[1]}$ and multiplying by $N$, we get
$(1+\frac{1}{2N})N\widehat{[1]}\leq Nk\widehat{v}+\widehat{[1]}.$
Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that
$sN[1]\leq(s+1)N[1]\leq s(Nkv+[1]).$
Since $sN[1]$ is compact, we can pick $v^{\prime}\in\operatorname{Cu}(C/I)$ such that
$v^{\prime}\ll v,\,\,\,\text{ and }\,\,\,sN[1]\leq s(Nkv^{\prime}+[1]).$
Using also that $v\leq[1]$ (since $kv\leq[1]$), we can apply [RR13, Lemma 2.3 (ii)] to find $e\in(C/I)_{+}$ such that $v^{\prime}\leq[e]\leq v$, and so
$k[e]\leq[1],\,\,\,\text{ and }\,\,\,sN[1]\leq s(Nk[e]+[1]).$
Separately, and using again that $[1]\in\operatorname{Cu}(C/I)$ satisfies condition (L’), we apply Lemma 4.8 to the compact element $[1]$ in $\operatorname{Cu}(C/I)$ and for the given $k$ and $\tfrac{1}{2N}$ to obtain $w\in\operatorname{Cu}(C/I)$ such that
$w\leq[1],\quad(1-\tfrac{1}{2N})k\widehat{w}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq kw.$
Multiplying by $N$, we get
$(N-\tfrac{1}{2})k\widehat{w}\leq N\widehat{[1]}.$
Using that $[1]$ is compact and $[1]\leq kw$, we find $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(C/I)$ such that
$w^{\prime\prime}\ll w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,[1]\leq kw^{\prime\prime}.$
First, applying Proposition 2.3(ii), we find a positive integer $t$ such that
$t(N-1)kw^{\prime}\leq(t+1)(N-1)kw^{\prime}\leq tN[1].$
Then, arguing as above, we find $f\in(C/I)_{+}$ with $w^{\prime\prime}\leq[f]\leq w^{\prime}$, and thus
$[f]\leq[1],\quad t(N-1)k[f]\leq tN[1],\,\,\,\text{ and }\,\,\,[1]\leq k[f].$
Let $\bar{e},\bar{f}\in C_{+}$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
$k[\bar{e}]\leq[1]+z,\quad sN[1]\leq s(Nk[\bar{e}]+[1])+z,$